# Exploring knowledge graph-based neural-symbolic system from application perspective

### v1

@@ 3.1.2 KG-based Q&A systems
, where $f_{\text{enc}}$ is the encoder extracting text features and $f_{\text{head}}$ predicts the relevance score of node $v$ to context $z$. Second is joint reasoning, where after forming the joint graph, the model uses a graph neural network with an attention mechanism to update node representations, iteratively updating the representations of the context and KG, thereby enabling reasoning. During this process, the Graph Attention Network (GAT)[125] dynamically adjusts the weights of information transmission between nodes, allowing each node to update its information based on the relevance score with its neighboring nodes.
@@ A.2 Graph neural networks
Graph neural networks (GNNs) are a class of neural networks designed to operate on graph-structured data, where entities (nodes) and their relationships (edges) are represented as a graph. Unlike traditional neural networks, which operate on grid-like data such as images or sequences, GNNs can learn and reason on highly irregular and interconnected data. At their core, GNNs iteratively update node representations by aggregating information from neighboring nodes, allowing them to capture complex dependencies and patterns within the graph structure. This process typically involves message passing between nodes, where each node aggregates information from its neighbors and updates its own representation accordingly.
@@ A.2.1 Graph convolutional networks
Graph convolutional networks (GCNs) are a type of graph neural network, which extend convolutional neural networks (CNNs) to handle irregular data domains such as graphs. In GCNs, convolutional operations are performed directly on the graph structure, allowing them to capture local and global information from neighboring nodes in a principled manner. The key idea behind GCNs is to aggregate feature information from a node’s neighborhood, typically represented by its immediate neighbors, and update the node’s representation accordingly. This process is repeated across multiple layers, allowing the network to learn hierarchical representations of the graph data.
@@ A.2.2 Graph attention networks
Graph Attention Networks (GATs) are a type of graph neural network that incorporates attention mechanisms to capture complex dependencies in graph-structured data. Unlike traditional graph convolutional networks (GCNs), which aggregate information from neighboring nodes uniformly, GATs dynamically assign attention weights to neighbors, allowing the model to focus on more relevant nodes during the aggregation process. This attention mechanism enables GATs to learn to weight the importance of different nodes based on their relevance to the target node, effectively capturing varying levels of importance in the graph structure. GATs have demonstrated superior performance compared to conventional GCNs, particularly in scenarios where nodes have varying degrees of importance or when dealing with large and heterogeneous graph structures.
A.2.3 Prototype graph neural networks
Prototype Graph Neural Networks (PGNNs) are a novel class of graph neural networks that leverage the concept of prototypes to enhance representation learning on graph-structured data. In PGNNs, prototypes are representative nodes or embeddings that capture the essence of different clusters or classes within the graph. Unlike traditional GNNs that learn representations based solely on local and global graph structure, PGNNs explicitly learn to identify and utilize prototype nodes, which serve as anchor points for the entire graph. These prototypes are learned during training through iterative refinement processes, where nodes are attracted to or repelled from them based on their similarity in feature space. By incorporating prototype-based learning, PGNNs can effectively capture the underlying structure and semantics of the graph, leading to more discriminative and interpretable representations.

### v2

@@ 3.1.2 KG-based Q&A systems
, where $f_{\text{enc}}$ is the encoder extracting text features and $f_{\text{head}}$ predicts the relevance score of node $v$ to context $z$. Second is joint reasoning, where after forming the joint graph, the model uses a graph neural network with an attention mechanism to update node representations, iteratively updating the representations of the context and KG, thereby enabling reasoning. In this process, the graph attention network (GAT)[116] dynamically adjusts the weights of information transfer between nodes so that each node can update its own information based on its relevance score with neighboring nodes.
@@ A.2 Graph neural networks
Graph Neural Networks (GNNs)[124] belong to a special class of neural networks that are designed to operate on data following the structure of a graph, in which there are entities or nodes, and relations or edges between them. That’s pretty uncommon, actually. Unlike conventional neural networks, which can learn and reason from grid-like data, GNNs are quite able to get pattern recognition within data that’s highly irregular and interconnected. GNNs iteratively update the representations of the nodes by allowing them to pool information from neighboring nodes, such that they can represent complex dependencies and patterns of the graph structure. This is the process, usually message passing between nodes, whereby each node aggregates information from its neighbors and changes its own representation based on the information.
@@ A.2.1 Graph convolutional networks
Graph convolutional networks (GCNs) [115]are, in fact, convolutional neural networks (CNNs) that generalize the classical CNNs’ convolutional operation from regular data domains to irregular ones, such as graphs. Being directly performed over the graph structure, convolutional operations in GCNs allow for the capturing of localized and global information in a principled manner over the corresponding adjacent nodes. Central to the idea of GCNs is the aggregation of feature information from the node’s neighborhood, normally understood to mean the immediate neighbors, followed by updating the node representation with the gathered representation, and repeating this process several times over multiple layers. This is being done several times over a few layers to progressively learn more complex hierarchical representations of the graph data.
@@ A.2.2 Graph attention networks
Graph Attention Networks (GATs)[116] are another type of graph neural network that applies attention mechanisms to allow the model to capture complex dependencies between different nodes in graph-structured data, within natural language processing. In traditional graph convolutional networks (GCNs), information from neighbors of nodes is uniformly aggregated, but the attention of GAT is dynamically computed for each pair of neighbor nodes and central node, which enabled our model to generalize the convolutional model by focusing more on the relevant neighbors of each node. This attention mechanism, in other words, allows the GATs to effectively learn the importance in light of their relevance to the target node, capturing differences in level that arise from the graph structure.